### v5

Abstract
Contemporary autonomous vehicle (AV) benchmarks have advanced techniques for training 3D detectors, particularly on large-scale multi-modal (LiDAR + RGB) data. While class labels naturally follow a long-tailed distribution in the real world, surprisingly, existing benchmarks only focus on a few common classes (e.g., pedestrian and car) and neglect many rare but crucial classes (e.g., emergency vehicle and stroller). However, AVs must reliably detect both common and rare classes for safe operation in the open world. We address this challenge by formally studying the problem of Long-Tailed 3D Detection (LT3D), which evaluates all annotated classes, including those in-the-tail. We address LT3D with hierarchical losses that promote feature sharing across classes, and introduce diagnostic metrics that award partial credit to “reasonable” mistakes with respect to the semantic hierarchy (e.g., mistaking a child for an adult). Further, we point out that rare-class accuracy is particularly improved via multi-modal late fusion (MMLF) of independently trained uni-modal LiDAR and RGB detectors. Such an MMLF framework allows us to leverage large-scale uni-modal datasets (with more examples for rare classes) to train better uni-modal detectors, unlike prevailing end-to-end trained multi-modal detectors that require paired multi-modal data. Finally, we examine three critical components of our simple MMLF approach from first principles: (1) whether to train 2D or 3D RGB detectors for fusion, (2) whether to match RGB and LiDAR detections in 3D or the projected 2D image plane, and (3) how to fuse matched detections. Extensive experiments reveal that (1) 2D RGB detectors achieve better recognition accuracy for rare classes than 3D RGB detectors, (2) matching on the 2D image plane mitigates depth estimation errors for better matching, and (3) score calibration and probabilistic fusion notably improves the final performance further. Our MMLF significantly outperforms prior work for LT3D, particularly improving on the six rarest classes from 12.8 to 20.0 mAP! Our code and models are available on our project page.
keywords:
Long-Tailed Distribution; 3D Detection; Multi-Modal Fusion; Late Fusion; Autonomous Vehicles; LiDAR; RGB.
1 Introduction
3D object detection is a key component in many robotics systems such as autonomous vehicles (AVs). To facilitate research in this space, the AV industry has released many large-scale 3D annotated multi-modal datasets [21, 7, 73, 64]. However, these datasets often benchmark on only a few common classes (e.g., car and pedestrian) and ignore rare classes (despite being annotated by some datasets [5]) like stroller and emergency vehicle (Fig. 1). In the real open world, safe navigation requires AVs to reliably detect rare classes such as debris and stroller [66, 75]. This motivates the study of Long-Tailed 3D Detection (LT3D), a problem that requires detecting objects from both common and rare classes.
Few      Medium      Many
Refer to caption Refer to caption
Figure 1: According to the histogram of per-class object counts (on the left), the nuScenes benchmark [5] focuses on the common classes in cyan (e.g., car and barrier) but ignores rare ones in red (e.g., stroller and debris). In fact, the benchmark creates a superclass pedestrian by grouping multiple classes in green, including the common class adult and several rare classes (e.g., child and police-officer); this complicates the analysis of detection performance as pedestrian performance is dominated by adult. Moreover, the ignored superclass pushable-pullable also contains diverse objects such as shopping-cart, dolly, luggage and trash-can as shown in the top row (on the right). We argue that AVs should also detect rare classes as they can affect AV behaviors. To comprehensively analyze performance, we report accuracy for three groups of classes (Many, Medium, and Few shown in the left panel) in the main paper and provide per-class performance in the appendix.
Status Quo. Among contemporary AV datasets, nuScenes [5] exhaustively annotates objects of various classes crucial for safe navigation (Fig. 1) and organizes them into a semantic hierarchy (Fig. 4). We are motivated to study LT3D by re-purposing all annotated classes in nuScenes because detecting rare classes is useful for downstream tasks such as motion planning. Importantly, LT3D is not simply solved by retraining state-of-the-art methods on both common and rare classes [46]. For example, CMT [80], a multi-modal transformer-based detector, achieves only 4.8 mAP on rare categories despite achieving 79.9 mAP on common classes (Table 1). We significantly extend our previous conference paper [46] which introduces LT3D by (1) studying whether to fuse 2D or 3D RGB detectors; (2) evaluating whether to fuse in 2D image plane or 3D LiDAR space; (3) investigating the impact of score calibration and probabilistic fusion; (4) rigorously comparing recently published end-to-end trained multimodal detectors.
Protocol. LT3D requires 3D localization and recognition of objects from each of the common (e.g., adult and car) and rare classes (e.g, child and stroller). Moreover, for safety-critical robots such as AVs, detecting but mis-classifying rare objects (e.g., mis-classifying a child as an adult) is preferable to failing to detect them at all. Therefore, we propose a new diagnostic metric to quantify the severity of classification mistakes in LT3D that exploits inter-class relationships w.r.t to the semantic hierarchy when awarding partial credit (Fig. 4). We use both the standard and proposed metrics to evaluate 3D detectors on all classes. Further, as prior works focus on only a few common classes, they miss opportunities to exploit this hierarchy in training. We propose hierarchical losses to promote feature sharing across common and rare classes, achieving improved performance.
Figure 2: We extensively explore the simple multi-modal late-fusion (MMLF) framework for LT3D by fusing RGB and LiDAR uni-modal detectors [46]. We rigorously examine three critical components within this framework (Fig. 3) and propose a simple method that fuses detections produced by a 2D RGB-detector (e.g., DINO [88]) and a 3D LiDAR-detector (e.g., CenterPoint [82]). Our method achieves 51.4 mAP on the nuScenes [5] LT3D benchmark, significantly improving over state-of-the-art detectors by 5.9% (Table 1).
Figure 3: We examine three key components in our multi-modal late-fusion (MMLF) of uni-modal RGB and LiDAR detectors from first principles: A. whether to train 2D or 3D RGB detectors, B. whether to match uni-modal detections on the 2D image plane or in the 3D bird’s-eye-view (BEV), and C. how to best fuse matched detections. Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining calibrated scores with Bayesian fusion yields state-of-the-art LT3D performance (Table 3 and Table 4).
Technical Insights. To address LT3D, we start by retraining state-of-the-art 3D detectors on all classes. Somewhat surprisingly, these detectors perform rather poorly on rare classes, e.g., CenterPoint [82] achieves only 0.1 AP on child and stroller (Table 5). We propose several algorithmic innovations to improve these results. First, we allow feature sharing across common and rare classes by training a single feature trunk, adding in hierarchical losses that ensure features will be useful for all classes (Table 4). Second, noting that LiDAR data is too impoverished for even humans to recognize certain rare objects that tend to be small (e.g., strollers), we propose a Multi-Modal Late-Fusion (MMLF) framework (Fig. 2) that fuses detections from a LiDAR-only detector (for precise 3D localization) and an RGB-only detector (for better recognition). Within MMLF, we first introduce a simple Multi-Modal Filtering (MMF) method, which post-processes predictions from 3D LiDAR-only (e.g., CenterPoint [82]) and 3D RGB-only (e.g., FCOS3D [76]) detectors, filtering away detections that are inconsistent across modalities. This improves LT3D performance by 3.2 mAP (Table 2). Next, we delve into this MMLF framework and study three critical design choices (Fig. 3): (A) whether to train a 2D or a 3D RGB detector for late-fusion, (B) whether to match detections in the 2D image plane or in 3D, and (C) how to optimally fuse detections. Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining score-calibrated predictions with probabilistic fusion yields state-of-the-art LT3D performance, significantly outperforming end-to-end trained multi-modal 3D detectors (Table 1).
Contributions. We present three major contributions. First, we formulate the problem of LT3D, emphasizing the detection of both common and rare classes in safety-critical applications like AVs. Second, we design LT3D’s benchmarking protocol and develop a diagnostic metric that awards partial credit depending on the severity of misclassifications (e.g., misclassifying child-vs-adult is less problematic than misclassifying child-vs-car). Third, we propose several architecture-agnostic approaches to address LT3D, including a simple multi-modal late-fusion (MMLF) strategy that generalizes across different RGB and LiDAR architectures (Appendix D). We conduct extensive experiments to ablate our design choices and demonstrate that our simple MMLF approach achieves state-of-the-art results on the nuScenes and Argoverse 2 LT3D benchmarks.
2 Related Work
3D Object Detection for AVs. Contemporary approaches to 3D object detection can be broadly classified as LiDAR-only, RGB-only, and sensor-fusion methods. Recent work in 3D detection is inspired by prior work in 2D detection [86, 33, 8]. LiDAR-based detectors like PointPillars [41], CBGS [84], and PVRCNN++ [63] adopt an SSD-like architecture [33] that regresses amodal bounding boxes from a bird’s-eye-view (BEV) feature map. More recently, CenterPoint [82] adopts a center-regression loss that is inspired by CenterNet [86]. Recent LiDAR-based detectors adopt the transformer architecture or attention modules that achieve further improvements [30, 34]. Despite significant progress, LiDAR-based detectors often produce many false positives because it is difficult to distinguish foreground objects from the background given sparse LiDAR returns. Monocular RGB-based methods have gained popularity in recent years due to increased interest in camera-only perception. FCOS3D [76] extends FCOS [67] by additionally regressing the size, depth, and rotation for each object. More recently, methods such as BEVDet and BEVFormer [25, 24, 43] construct a BEV feature-map by estimating the per-pixel depth of each image feature [47]. PolarFormer [31] introduces a polar-coordinate transformation that improves near-field detection. Many of these state-of-the-art 3D RGB detectors are commonly pre-trained on large external datasets like DDAD [17]. Monocular RGB detectors accurately classify objects but struggle to estimate depth, particularly for far-field detections [20]. Despite recent advances in LiDAR and RGB 3D detectors, we find that multi-modal fusion is essential for LT3D (detailed next). Importantly, using both RGB (for better recognition) and LiDAR (for better 3D localization) helps detect rare classes. Our work focuses on multi-modal late-fusion (Fig. 2) and studies how to effectively fuse RGB and LiDAR uni-modal detectors for LT3D.
Multi-modal 3D Detection. Multi-modal fusion for 3D detection is an active field of exploration [2, 40, 81]. Existing methods can be categorized as input-fusion, feature-fusion, and late-fusion. Input-fusion typically augments LiDAR points using image-level features. For example, PointPainting [69] projects LiDAR points onto the output mask of a semantic segmentation model and appends corresponding class scores to each point. MVP [83] densifies regions of LiDAR sweeps that correspond with objects in semantic segmentation masks. In contrast, Frustum PointNets [50] leverage 2D RGB detections to localize objects within the box frustum using PointNets [51].
Recent works show that feature-fusion is more effective than input-fusion. [77] fuses global image and point-cloud features prior to detection; [29, 11] fuse LiDAR and RGB features at multiple scales. TransFusion [2] and BEVFusion [40] fuse features in the BEV space using multi-headed attention. Despite the success of transformers for detecting common objects, [46] finds that TransFusion struggles to detect rare classes, and posits that the transformer architecture, as adopted in TransFusion and BEVFusion, suffers from limited training data (particularly for classes in the long tail). For transformers to work well in practice, they should be trained on diverse, large-scale datasets [12, 59]. It is also worth noting that end-to-end trained multi-modal detectors require paired multi-modal data for training, increasing the cost of data collection and modality alignment. In contrast, our work focuses on multi-modal late-fusion (MMLF), a framework that ensembles uni-modal detectors, which do not require aligned RGB-LiDAR paired training data. CLOCs [49] is a late-fusion method that learns a separate network to fuse RGB and LiDAR detections, showing promising results for 3D detection. However, CLOCS only performs late fusion on matched predictions with semantic agreement and does not fix misclassifications. We find that handling such misclassifications is critical for improving rare class performance. We delve into this simple MMLF framework, study three crucial design choices, and present the final method that significantly outperforms the state-of-the-art approaches for LT3D.
Long-Tailed Perception. AV datasets follow a long-tailed class distribution: a few classes like car and pedestrian are dominant, while others like stroller and debris are rarely seen. In fact, Long-Tailed Perception (LTP) is not a unique problem in the AV domain [54] but a long-standing problem in the literature [37] and has been widely studied through the lens of image classification, aiming for high accuracy averaged across imbalanced classes [37, 85, 1]. Existing LTP methods propose reweighting losses [6, 32, 26, 89], rebalancing data sampling [14, 4, 27], balancing gradients computed from imbalanced classes [65], and balancing network weights [1]. Others study LTP through the lens of 2D object detection [19]. Compared to long-tailed image classification, long-tailed 3D detection provides unique opportunities and challenges as sensors such as LiDAR directly provide geometric and ego-motion cues that are difficult to extract from 2D images. Further, 2D detectors must detect objects of different scales due to perspective image projection, dramatically increasing the complexity of the output space (e.g., requiring more anchor boxes). In contrast, 3D objects do not exhibit as much scale variation, but far-away objects tend to have sparse LiDAR returns [20, 48], imposing different challenges. Finally, 3D detectors often use class-aware heads (i.e. each class has its own binary classifier) while 2D long-tail recognition approaches typically use shared softmax heads.
Recently, CBGS [84] explicitly addresses rare-class 3D detection by up-sampling LiDAR-sweeps with instances of rare classes and pasting instances of rare objects copied from different scenes. Although this works well for improving detection of infrequently-seen classes (e.g. classes with medium number of examples like bicycle and construction vehicle), it does not provide significant improvement for classes with only a few examples like debris and stroller [46]. Additionally, rare classes, such as child and stroller, are typically small in size and have a limited number of LiDAR returns. As a result, LiDAR-only detectors struggle to accurately recognize these rare classes. In this work, we address LT3D by multi-modal late-fusion, which ensembles RGB and LiDAR uni-modal detectors.
3 LT3D Benchmarking Protocol
[…]
Mean Average Precision (mAP) is a well-established metric for object detection [15, 21, 36]. For 3D detection on LiDAR sweeps, a true positive is defined as a detection that has a center distance within a distance threshold to a ground-truth annotation [5]. mAP computes the mean of AP over classes, where per-class AP is the averaged area under the precision-recall curves with distance thresholds of [0.5, 1, 2, 4] meters.
Hierarchical Mean Average Precision (mAPH). For safety-critical applications, it is more desirable to localize-but-misclassify an object than fail to detect this object, e.g., detecting but misclassifying a child as adult is better than not detecting this child. Therefore, we propose hierarchical mean average precision (mAPH), which considers such semantic relationships across classes to award partial credit. We use APH as a supplementary diagnostic tool to analyze how LT3D detectors make mistakes. To encode inter-class relationships, we leverage the semantic hierarchy (Fig. 4) such as the one defined by nuScenes [5]. We derive partial credit as a function of semantic similarity using the least common ancestor (LCA) distance metric. LCA has been proposed for image classification [53, 62] but not for object detection. Extending this metric to detection is non-trivial as we must jointly evaluate semantic and spatial overlap. Below, we explain how to compute APH for a class $C$.
LCA=0: Consider the predictions and ground-truth boxes for $C$. Label the set of predictions that overlap with ground-truth boxes for $C$ as true positives. Other predictions are false positives. This is identical to the standard mAP metric.
LCA=1: Consider the predictions for $C$, and ground-truth boxes for $C$ and its all sibling classes with LCA distances to $C$ of 1. Label the predictions that overlap a ground-truth box of $C$ as a true positive. Label the predictions that overlap sibling classes as ignored [36]. All other predictions for $C$ are false positives.
LCA=2: Consider the predictions for $C$ and ground-truth boxes for $C$ and all its sibling classes with LCA distances to $C$ no greater than $2$. For nuScenes, this includes all classes. Label the set of predictions that overlap ground-truth boxes for $C$ as true positives. Label the set of predictions that overlap other classes as ignored. All other predictions for $C$ are false positives.
Figure 4: The nuScenes dataset defines a hierarchy for the annotated classes (Fig. 1). We mark common classes in white and rare classes in gold. The standard nuScenes benchmark makes two choices for dealing with rare classes: (1) ignoring them (e.g., stroller and pushable-pullable), or (2) grouping them into super classes (e.g., adult, child, construction-worker, police-officer are grouped as pedestrian). As the pedestrian class is dominated by adult (Fig. 1), the standard benchmarking protocol masks the challenge of detecting rare classes like child and police-officer. We explicitly incorporate this hierarchy to evaluate LT3D performance via the proposed mAPH metric described in Sec. 3.
4 Our Proposed Approach
To address LT3D, we first retrain well-established detectors on all long-tailed classes. We consider the following detectors:
  • LiDAR-based 3D Detectors: CenterPoint [82], PointPillars [41], TransFusion-L [2], BEVFusion-L [40], CMT-L [80];
  • RGB-based 3D Detectors: FCOS3D [76], PolarFormer [31], BEVFormer [43];
  • RGB-based 2D Detectors: YOLOV7 [70], and DINO [88];
  • End-to-End Multi-modal 3D Detectors: TransFusion [2], BEVFusion [40], CMT [80], DeepInteraction [79], IS-Fusion [81], and EA-LSS [28].
In Section 4.1 and 4.2, we introduce two modifications that consistently improve LT3D performance. In Section 4.3, we present a simple yet effective late-fusion method.
4.1 Grouping-Free Detector Head
Extending existing 3D detectors to train with more classes is surprisingly challenging. Many contemporary networks use a multi-head architecture that groups classes of similar size and shape to facilitate efficient feature sharing [84, 82]. For example, CenterPoint [82] groups pedestrian and traffic-cone since these objects are both tall and skinny. This multi-head grouping design may not work for super-classes like pushable-pullable and debris that contain diverse objects of different sizes and shapes. Moreover, in the multi-head design, each head is a group-specific detector that consists of several layers with lots of parameters. This makes multi-head grouping difficult to scale for a large number of classes. To address these issues, we treat each class as its own group to avoid hand-crafted grouping heuristics and design a group-free strategy, in which each class has only one linear layer as its detector (Fig. 5) and all classes share a single detector head. This design significantly reduces the number of parameters and allows learning the shared feature backbone collaboratively with all classes, effectively mitigating overfitting to rare classes. Adding a new class is as simple as adding a single linear layer to the detector head. Importantly, we show that our grouping-free detector outperforms conventional grouping-based methods in Appendix F.
Figure 5: We leverage the semantic hierarchy defined in the nuScenes dataset (Fig. 4) to train LT3D detectors by predicting class labels at multiple levels of the hierarchy for an object: its fine-grained label (e.g., child), its coarse class (e.g., pedestrian), and the root-level class object. This means that the final vocabulary of classes is no longer mutually exclusive, so we use a sigmoid focal loss that learns separate spatial heatmaps for each class.
4.2 Training with a Semantic Hierarchy
The nuScenes dataset defines a semantic hierarchy (Fig. 4) that groups semantically similar classes under coarse-grained categories [5]. We leverage this hierarchy during training. Specifically, we train detectors to predict three labels for each object: its fine-grained class (e.g., child), its coarse class (e.g., pedestrian), and the root class object. We adopt a grouping-free detector head (Fig. 5) that outputs separate “multitask” heatmaps for each class, and use a per-class sigmoid focal loss [35] rather than multi-class cross-entropy loss to avoid normalizing class probabilities across fine-grained and super-classes. It is worth noting that our approach does not explicitly enforce a tree-like hierarchy, and can be applied to more complex label relationships [39]. Crucially, adding a vehicle heatmap does not directly interfere with the car heatmap. However, this might produce repeated detections on the same test object (e.g. a single ground-truth car may be detected as a car, vehicle, and object). We address that by simply ignoring coarse detections at test time. We explore alternatives in Appendix G and conclude that they achieve similar LT3D performance. Perhaps surprisingly, this training method improves detection performance not only for rare classes, but also for common classes (Table 4).
4.3 Multi-Modal Filtering for Detection Fusion
Rare-class instances are often small and can be challenging to recognize from sparse (LiDAR) geometry alone: even humans struggle to find strollers in LiDAR point clouds. This suggests that one can leverage multi-modal cues to improve LT3D. First, we find that, although LiDAR-based detectors are widely adopted for 3D detection, they produce many high-scoring false positives (FPs) for rare classes due to misclassification. We focus on removing such FPs. To this end, we use an RGB-based 3D detector (e.g., FCOS3D [76]) to filter out high-scoring false-positive LiDAR detections by leveraging two insights: (1) LiDAR-based 3D-detectors can achieve high recall and precise 3D localizations for TPs, and (2) RGB-based 3D-detections are accurate w.r.t recognition although their 3D localization is poor. Fig. 6 demonstrates our Multi-Modal Filtering (MMF) strategy. For each RGB detection, we keep LiDAR detections within a radius of $m$ meters and remove all the other LiDAR detections. We denote this method as MMF($D_{L}$, $D_{R}$), where DL and $D_{R}$ are any 3D LiDAR detector and 3D RGB detector, respectively. Table 2 demonstrates that MMF greatly improves LT3D.
LiDAR detections RGB detections Filtered results
Figure 6: The proposed method Multi-Modal Filtering (MMF) effectively removes high-scoring false-positive LiDAR detections. The green boxes are ground-truth strollers, while the blue boxes are stroller detections from the LiDAR-based detector CenterPoint [82] (left) and RGB-based detector FCOS3D [76] (mid). The final filtered result removes LiDAR detections not within $m$ meters of any RGB detection, shown in the red region, and keeps all other LiDAR detections, shown in the white region (right).
Limitations of MMF. Despite the effectiveness of multi-modal filtering, our approach is sensitive to the classification accuracy of FCOS3D when matching LiDAR and RGB-based detections, leading to many correctly localized but misclassified detections. We use a confusion matrix to further analyze the misclassifications within each superclass, as shown in Fig. 7. We explain how to compute a confusion matrix for the detection task: For each superclass, we construct a confusion matrix, in which the entry $(i,j)$ indicates the misclassification rate of class-$i$ objects as class-$j$. Specifically, given a class $i$, we find its predictions that match ground-truth boxes within a 2m center-distance of class-$i$ and all its sibling classes (LCA=1, within the corresponding superclass); we ignore all unmatched detections. This allows us to count the mis-classifications of class-$i$ objects into class-$j$. We normalize these counts to produces misclassification rates. We find that rare classes are most often confused by the dominant class in each superclass: Vehicle is dominated by car, Pedestrian is dominated by adult, and Movable is dominated by barrier. We posit that this can be addressed with more accurate RGB-based classifiers, motivating our study of multi-modal late-fusion (MMLF) from first principles, presented next.
Vehicle     Pedestrian     Movable
Refer to caption Refer to caption Refer to caption
Figure 7: We analyze our multi-modal filtering (MMF) approach. Rare classes are most often confused by the common class (in blue) in each superclass: (left) Vehicle is dominated by car, (mid) Pedestrian is dominated by adult, and (right) Movable is dominated by barrier. We find that class confusions are reasonable. Car is often mistaken for truck. Similarly, truck, construction-vehicle and emergency-vehicle are most often mistaken for car. Bicycle and motorcycle are sometimes misclassified as car, presumably because they are sometimes spatially close to cars. Adults have similar appearance to police-officer and construction-worker, and they are often co-localized with child and stroller.
5 Multi-Modal Late-Fusion for LT3D
Motivated by the effectiveness of our frustratingly simple multi-modal filtering (MMF) approach, we investigate this multi-modal late-fusion (MMLF) strategy further. We explore three questions from first principles (Fig. 3): (1) how to effectively incorporate RGB information, (2) how to match RGB and LiDAR detections, and (3) how to fuse them.
5.1 How Do We Incorporate RGB Information?
Although LiDAR offers accurate localization, LiDAR-only detectors struggle to recognize objects using sparse LiDAR alone. RGB images provide complementary information that is essential for identifying objects and disambiguating those that are geometrically similar in point clouds but semantically different in images, e.g., construction worker vs. police officer. In the last subsection, we proposed multimodal filtering (MMF) (Fig. 6), a simple late-fusion strategy that ensembles a 3D RGB detector and a 3D LiDAR detector and yields remarkable improvements for LT3D (Table 2). Motivated by the late-fusion strategy, we further explore fusing a 2D RGB detector and a 3D LiDAR detector. Surprisingly, this yields more significantly better LT3D performance. We present insights on why using a 2D detector to incorporate RGB information is better and ablate the impact of using 2D vs. 3D RGB detectors for late-fusion in Table 3.
Figure 8: Our multi-modal late-fusion (MMLF) strategy takes 3D LiDAR and 2D RGB detections as input (Fig. 3A), matches the detections on the image plane (Fig. 3B), and fuses matched ones with score calibration and probabilistic fusion to produce the final 3D detections (Fig. 3C).
2D RGB detectors are more mature. 2D object detection is a fundamental problem in computer vision [16, 36, 57] that has matured in recent years and model trade-offs are well understood [57, 33, 52, 35]. Our work considers two well-established 2D RGB detectors, YOLOV7 [70] and DINO [88]. YOLOV7 is a real-time detector that identifies a number of training techniques that nearly doubles the inference efficiency over prior work without sacrificing performance. DINO is a transformer-based detector that improves upon DETR [8] using denoising anchor boxes. As 2D detectors do not estimate 3D attributes (e.g., depth and rotation), understanding how to best leverage them in the context of LT3D is a key challenge (explored in Section 5.2).
2D RGB detectors can be trained with more diverse data. Training 2D RGB detectors only requires 2D bounding box annotations, which are significantly cheaper to collect than 3D cuboids used for training 3D RGB detectors [76, 31]. Since annotating 3D amodal cuboids is both expensive and non-trivial (compared to bounding-box annotations for 2D detection), datasets for monocular 3D RGB detection are considerably smaller and less diverse than their 2D detection counterparts. For example, nuScenes [5] (published in 2020) annotates 144K RGB images of 23 classes using 3D cuboids, while COCO [36] (an early 2D detection dataset published in 2014) annotates 330K images of 80 classes using 2D bounding boxes. This allows us to pretrain 2D RGB detectors on significantly larger, more diverse, publicly available datasets [44, 87, 71, 78, 55]. We demonstrate in Fig. 11 that leveraging existing 2D detection datasets helps train stronger 2D detectors, further improving LT3D performance.
5.2 How Do We Match Uni-Modal Detections?
Finding correspondence between two sets of RGB and LiDAR uni-modal detections is an essential step in the late-fusion framework (Fig. 3B). Our previous MMF method matches 3D RGB and 3D LiDAR detections using center distance in the 3D space (Fig. 6). However, precisely matching them in the 3D space is error-prone due to depth estimation errors from the 3D RGB detector. Hence, we opt to match 2D RGB (without the need to estimate depth in detection) and 3D LiDAR detections. Prior work attempts to inflate 2D detections to 3D for matching by using LiDAR points [72], but we find that matching inflated 2D detections also yields noisy 3D boxes (due to outlier LiDAR points from the background) and reduces overall match quality. In contrast, we match uni-modal detections by projecting 3D LiDAR detections onto the 2D image plane, avoiding additional noise due to imprecise depth estimates. We ablate the impact of matching in 3D versus on the 2D image plane in Table 3, and present our 2D matching algorithm below.
Spatial matching on the 2D image plane. Using the available sensor extrinstics, we project 3D LiDAR detections onto the 2D image plane. We then use the IoU metric to determine overlap between projected LiDAR and 2D RGB detections. A 2D RGB detection and a projected 3D LiDAR detection are considered a match if their IoU is greater than a threshold. Although conceptually simple, this matching method works significantly better than using center distance to match detections in 3D (cf. bottom two rows in Table 3). Spatially matching uni-modal detections using 2D IoU yields three categories of detections: matched detections, unmatched RGB detections (that do not have corresponding LiDAR detections), and unmatched LiDAR detections (that do not have corresponding RGB detections). We handle unmatched detections below and show how to fuse matched ones in the next subsection.
Handling unmatched detections. We remove unmatched 2D RGB detections, positing that any unmatched RGB detections are likely to be false positives given that LiDAR detectors tend to yield high recall [46] (Appendix M). For unmatched 3D LiDAR detections, we down-weight their detection confidence scores by $w$ (inspired by SoftNMS [3]). We optimize for $w$ by grid search on a validation set and set $w=0.4$.
5.3 How Do We Fuse Matched Uni-Modal Detections?
As illustrated by Fig. 3C and Fig. 9, detections may match spatially but not semantically. To address this, we propose a semantic matching heuristic. Given a pair of spatially matched RGB and LiDAR detections, we consider two cases: matched detections with semantic disagreement (e.g., RGB and LiDAR predict different classes), and matched detections with semantic agreement (e.g., RGB and LiDAR predict the same class).
Addressing semantic disagreement between modalities. If the two modalities predict different semantic classes, we use the confidence score (which is calibrated as explained below) and class label of the RGB-based detection, and the 3D box from the LiDAR-based detection. Intuitively, RGB detectors can predict semantics more reliably from high resolution images than LiDAR-only detectors. This helps correct misclassifications of geometrically similar but semantically different objects produced by the 3D LiDAR detector, as shown in Fig. 9. To note, prior late-fusion methods like CLOCS [49] and our MMF [46] only fuse matched predictions with semantic agreement and do not correct misclassifications. In contrast, ours handles misclassifications that significantly improves rare class performance (Table 1).
Fusing semantically-agreed detections. If both modalities predict the same semantic class, we perform score fusion and probablistic fusion [9] as described below. Note that the confidence scores of RGB and LiDAR detections are not directly comparable: LiDAR-based detectors are often under-confident as it is difficult to distinguish foreground-vs-background using sparse LiDAR alone. Therefore, score calibration is crucial for fusion. Below, we explore score calibration of RGB ($x_{\text{RGB}}$) and LiDAR ($x_{\text{LiDAR}}$) detections.
[…]
Probabilistic fusion. We assume independent class prior $p(c)$ and conditional independence given the class label $c$ [9], i.e., $p(x_{\text{RGB}},x_{\text{LiDAR}}|c)=p(x_{\text{RGB}}|c)p(x_{\text{LiDAR}}|c)$. We compute the final score as
$\small p(c|x_{\text{RGB}},x_{\text{LiDAR}})$ $=\frac{p(x_{\text{RGB}},x_{\text{LiDAR}}|c)p(c)}{p(x_{\text{RGB}},x_{\text{LiDAR}})}$
$\propto\ \ p(x_{\text{RGB}},x_{\text{LiDAR}}|c){p(c)}$
$\propto\ \ p(x_{\text{RGB}}|c)p(x_{\text{LiDAR}}|c){p(c)}$
$\propto\ \ \frac{p(c|x_{\text{RGB}})p(c|x_{\text{LiDAR}})}{{p(c)}}$
where $p(c|x_{\text{RGB}})$ and $p(c|x_{\text{LiDAR}})$ are the posteriors after calibration. Unlike assuming balanced class prior distributions in [9], we tune $p(c)$ in a greedy fashion for classes sorted by their cardinality. Similar to temperature tuning in score calibration, probabilistic fusion does not increase inference time (Appendix E).
Method Modality All Many Medium Few
FCOS3D [76] (ICCVW’21) C 20.9 39.0 23.3 2.9
BEVFormer [43] (ECCV’22) C 27.3 52.3 31.6 1.4
PolarFormer [31] (AAAI’23) C 28.0 54.0 31.6 2.2
CenterPoint [82] (CVPR’21) L 40.4 77.1 45.1 4.3
TransFusion-L [2] (CVPR’22) L 38.5 68.5 42.8 8.4
BEVFusion-L [40] (ICRA’23) L 42.5 72.5 48.0 10.6
CMT-L [80] (ICCV’23) L 34.7 73.4 35.9 1.1
CLOCS [49] (IROS’20) C+L 40.0 68.2 45.7 10.0
TransFusion [2] (CVPR’22) C+L 39.8 73.9 41.2 9.8
DeepInteraction [79] (NeurIPS ’22) C+L 43.7 76.2 51.1 7.9
BEVFusion [40] (ICRA’23) C+L 45.5 75.5 52.0 12.8
CMT [80] (ICCV’23) C+L 44.4 79.9 53.0 4.8
EA-LSS [28] (arXiv’23) C+L 44.6 79.5 52.6 6.2
IS-Fusion [81] (CVPR’24) C+L 44.1 77.9 59.4 3.8
CenterPoint [82] + RCNN [18] C+L 34.0 64.8 37.5 4.3
MMF(CenterPoint, FCOS3D) C+L 43.6 77.1 49.0 9.4
MMLF(CenterPoint, DINO) C+L 51.4 77.9 59.4 20.0
Table 1: Benchmarking results on nuScenes. We denote the RGB and LiDAR modalities by C and L, respectively. Our simple late-fusion method MMF, which uses detections of the 3D RGB-based detector FCOS3D to filter detections of the 3D LiDAR-based detector CenterPoint, rivals prior methods. Furthermore, our final method MMLF, which fuses 3D LiDAR and 2D RGB detections (from CenterPoint and DINO, respectively) with score calibration and probabilistic fusion, performs the best averaged across all common and rare classes. Notably, it outperforms end-to-end multi-modal methods such as BEVFusion [40], DeepInteraction [79], and CMT [80]. Importantly, our MMLF nearly doubles the detection performance achieved by prior work on classes with few examples!
6 Experiments
We conduct extensive experiments to better understand the LT3D problem and gain insights by validating our techniques described in Sec. 4. We aim to answer the following questions: [^1] Answers: yes, yes, yes, yes, yes.
  1. 1.
    Are rare classes more difficult to detect than common classes?
  2. 2.
    Are objects from rare classes sufficiently localized but mis-classified?
  3. 3.
    Does training with the semantic hierarchy improve detection performance for LT3D?
  4. 4.
    Does our multimodal late-fusion strategy resoundingly outperform previous methods?
  5. 5.
    Does our multi-modal late-fusion help detect rare classes?
We compare our Multi-Modal Filter (MMF) and Multi-Modal Late-Fusion (MMLF) approaches (Fig. 6 and 8) with prior works and present a detailed ablation study to further addresses the three motivating questions in Fig. 3. Regarding notation, MMF(X, Y) and MMLF(X, Y) mean that they fuse detections from the LiDAR-based detector X and the RGB-based detector Y. Our approach improves over prior works by 5.9 mAP, notably improving by 7.2 mAP on rare classes (Table 1). We provide implementation details in Appendix A.
Datasets. To explore LT3D, we use the well-established nuScenes [5] and Argoverse 2.0 (AV2) [73] datasets. Other AV datasets such as KITTI [21] and Waymo [64] do not support the study of LT3D as they annotate only three common classes. In contrast, nuScenes and AV2 have 18 and 26 fine-grained classes, respectively, which follow long-tailed distributions. To quantify the long-tail, we compute the imbalance factor (IF), defined as the ratio between the numbers of annotations of the most and the least common classes [10]: nuScenes and AV2 have an IF=1,670 and 2,500 respectively, which are significantly more imbalanced than existing long-tail image recognition benchmarks, e.g., iNaturalist (IF=500) [68] and ImageNet-LT (IF=1,000) [38]. nuScenes arranges classes in a semantic hierarchy (Fig. 4); AV2 does not provide a semantic hierarchy but we construct one based on the nuScenes’ hierarchy (refer to Appendix B for details). Following prior work, we use the official training set for training and the validation set for benchmarking. As our primary conclusions hold for both datasets, we report results on nuScenes in the main paper and provide results on AV2 in Appendix B.
Figure 9: Three examples demonstrate how our multi-modal late-fusion (MMLF) approach improves LT3D by fusing 2D RGB detections (from DINO [88]) and 3D LiDAR detections (from CenterPoint [82]). In all examples, MMLF correctly relabels detections which are geometrically similar (w.r.t size and shape) in LiDAR but are visually distinct in RGB, such as bus-vs-truck, adult-vs-stroller, and adult-vs-child. Refer to Appendix O for 3D visualizations.
Method Modality All Many Medium Few
FCOS3D C 20.9 39.0 23.3 2.9
PointPillars (PP) L 30.0 64.2 28.4 3.4
 + Hierarchy Loss L 31.2 66.4 30.4 2.9
 + MMF(PP, FCOS3D) C+L 35.8 66.2 41.0 4.4
CenterPoint (CP) L 39.2 76.4 43.1 3.5
 + Hierarchy Loss L 40.4 77.1 45.1 4.3
 + MMF(CP, FCOS3D) C+L 43.6 77.1 49.0 9.4
[…]
6.1 Benchmarking Results
Table 1 compares our MMF and MMLF with prior work on nuScenes. Fig. 9 displays qualitative results. We adapt existing methods (which were previously trained on 10 classes in nuScenes) for LT3D by retraining them on all 18 classes. We provide breakdown analysis per class in Appendix H and show results on the standard 10-class benchmark in Appendix K.
CenterPoint [82], a popular LiDAR-only 3D detector, is unable to detect rare objects, achieving only 4.3 mAP on classes with few examples. This is expected as it is difficult to identify rare objects from sparse LiDAR points alone. Yet, it is unexpected that the transformer-based 3D LiDAR detector BEVFusion-L performs considerably better on few classes, achieving 10.6 mAP, but it underperforms CenterPoint by 4.6 mAP on many classes. We posit that the limited training data in-the-tail and class imbalance make it difficult to learn generalizable features, preventing robust LT3D performance (Appendix D). In contrast, BEVFusion [43], which is an end-to-end trained multi-modal method, performs 3.0 mAP better than the LiDAR-only variant (BEVFusion-L), confirming the benefit of using both RGB and LiDAR for LT3D.
 Method Fusion in 3D (MMF) Fusion in 2D (MMLF)
All Many Medium Few All Many Medium Few
   CenterPoint 40.4 77.1 45.1 4.3 40.4 77.1 45.1 4.3
+ FCOS3D 42.9 76.6 48.7 8.1 42.6 75.0 49.4 7.7
+ BEVFormer 43.2 76.9 50.8 6.3 42.8 75.2 51.4 5.7
+ PolarFormer 42.8 76.8 50.0 6.1 42.6 75.1 51.1 5.6
+ YOLOV7 40.1 76.1 43.8 5.8 45.7 77.1 52.8 11.2
+ DINO 40.3 76.2 44.1 5.9 49.5 77.4 57.7 16.7
Table 3: Fusing uni-modal detections in 3D vs. on the 2D image plane. We evaluate the impact of fusing 3D LiDAR detections (from CenterPoint trained with our hierarchy loss) with 2D RGB vs. and 3D RGB detections, in 3D vs. on the image plane. We match and filter detections in 3D using center distance (MMF), and match and filter detections in the 2D image plane using IoU (MMLF). When fusing in 3D, we follow [72] that inflates 2D detections to 3D using LiDAR points within their box frustums. When fusing on the 2D image plane, we project 3D detections to the image plane using provided sensor extrinsics. When fusing 3D RGB detections, similar results are achieved regardless matching in 3D or on the 2D image plane. Unsurprisingly, inflating 2D RGB detections for matching in 3D performs worse than matching 3D RGB detections in 3D. Yet, late-fusion between LiDAR-based 3D detections and RGB-based 2D detections in the image plane (bottom right panel) significantly improves performance for classes with medium and few examples by >>10 mAP. Refer to Appendix I for more results.
Method $\Delta$ All Many Medium Few
CenterPoint [82] 39.2 76.4 43.1 3.5
 + Hierarchy Loss +1.2 40.4 77.1 45.1 4.3
 + MMLF w/ DINO +7.5 47.9 77.1 55.8 14.4
 + External Data +1.9 49.8 77.1 57.1 18.6
 + Score Calibration +0.7 50.5 77.8 58.2 18.7
 + Probabilistic Fusion +0.9 51.4 77.9 59.4 20.0
Table 4: Adding our developed components progressively improves LT3D performance, confirming that (1) 2D RGB detectors are better suited for our MMLF, (2) matching projected 3D LiDAR detections on the 2D image plane outperforms matching 2D RGB detections inflated to 3D, and (3) score calibration prior and probabilistic fusion improves performance. A majority of performance improvements is attributed to fusing detections of the 2D RGB detector DINO on the 2D image plane. Moreover, training with external data notably improves rare class accuracy by 4.2 mAP.
Next, we implement another simple baseline, termed “CenterPoint + RCNN”, that trains a region-based CNN (RCNN) [18] classifier on cropped regions corresponding to projected 3D detections. Notably, it underperforms CenterPoint by 6.4 mAP, likely because it learns classifiers on cropped regions that does not exploit contextual information useful for recognition. This suggests that late-fusion cannot be simply solved with a 3D LiDAR detector and a strong 2D classifier. Our multimodal filtering method MMF keeps CenterPoint detections that are close to monocular 3D RGB detections produced by FCOS3D in 3D and discards all other LiDAR predictions. This simple baseline achieves 9.4 mAP on classes with few examples. Lastly, we re-implement CLOCS [49] by fusing detections from DINO and CenterPoint respectively. Note that CLOCS only fuses predictions from the same class, which prevents re-labeling misclassified LiDAR detections. However, CLOCs performs worse than our MMF baseline. By carefully considering design choices outlined in Fig. 3, MMLF improves over MMF by 7.8 mAP!
6.2 Ablation Study
We conduct extensive ablations to validate our proposed MMLF approach. In appendices, we provide additional results on the impact of calibration noise (Appendix N), detector recall (Appendix M) and performance breakdowns by object visibility and distance from ego-vehicle (Appendix L).
Method $mAP_{H}$ Car Adult Truck CV Bicycle MC Child CW Stroller $P$
CenterPoint [82] LCA=0 82.4 81.2 49.4 19.7 33.6 48.9 0.1 14.2 0.1 21.7
(original) LCA=1 83.9 82.0 58.7 20.5 35.2 50.5 0.1 18.3 0.1 22.0
LCA=2 84.0 82.4 58.8 20.7 36.4 51.0 0.1 19.5 0.1 22.6
CenterPoint (w/ our group-free head) LCA=0 88.1 86.3 62.7 24.5 48.5 62.8 0.1 22.2 4.3 32.7
LCA=1 89.0 87.1 71.6 26.7 50.2 64.7 0.1 29.4 4.5 32.9
LCA=2 89.1 87.5 71.7 26.8 51.1 65.2 0.1 30.5 4.8 33.4
CenterPoint (w/ our group-free head) LCA=0 88.6 86.9 63.4 25.7 50.2 63.2 0.1 25.3 8.7 36.8
+ Hierarchy Loss LCA=1 89.5 87.6 72.4 27.5 52.2 65.2 0.1 32.4 9.4 37.0
LCA=2 89.6 88.0 72.5 27.7 53.2 65.7 0.1 34.0 9.8 37.6
CenterPoint (w/ our group-free head) LCA=0 86.3 87.7 60.6 35.3 70.0 75.9 8.8 55.9 37.7 58.1
+ Hier. + MMLF LCA=1 86.8 88.3 68.5 37.3 70.4 77.1 16.2 66.0 51.5 58.2
LCA=2 86.9 88.6 68.6 37.7 70.9 77.4 16.3 69.0 52.4 58.9
CMT [80] LCA=0 88.6 87.7 65.2 36.9 66.7 76.3 4.7 34.4 0.9 34.1
(RGB + LiDAR) LCA=1 89.1 88.3 73.3 38.8 67.6 77.2 7.3 50.6 1.2 34.6
LCA=2 89.1 88.6 73.4 39.0 68.3 77.7 7.7 52.6 1.3 35.6
CMT-L (LiDAR-only) LCA=0 87.1 87.3 62.7 35.3 72.0 76.1 11.9 53.9 29.6 56.4
+ MMLF LCA=1 87.6 87.8 69.6 36.4 72.3 76.8 16.7 63.4 50.6 56.6
LCA=2 87.6 88.1 69.7 36.7 72.7 77.3 16.9 66.3 53.2 57.3
Table 5: Diagnosis using the mAPH metric on selected classes. We analyze the LiDAR-based 3D detector CenterPoint [82] and mutli-modal detector CMT [80], with our proposed group-free detector head, hierarchy loss, and MMLF (using the 2D RGB detector DINO [88]). Comparing the rows of LCA=0 for CenterPoint and CMT, we see that MMLF significantly improves on rare classes (in blue), e.g., construction-vehicle (CV), bicycle, motorcycle (MC), construction-worker (CW), stroller, and pushable-pullable ($P$). Moreover, performance increases significantly from LCA=0 to LCA=1 compared against LCA=1 to LCA=2, suggesting that objects from rare classes are often detected but misclassified as sibling classes.
6.2.1 Analysis on Training with Semantic Hierarchy
We validate the benefits of training with semantic hierarchy by modifying LiDAR-based detectors to jointly predict class labels at different levels of the hierarchy. For example, we modify detectors PointPillars [41] and CenterPoint [82] to additionally classify stroller as pedestrian and object. As the hierarchy naturally groups classes based on shared attributes, training with it helps learn complementary features across levels that are shared across common and rare classes. Table 2 highlights that this approach (“+ Hierarchy Loss”) achieves improvements on all class by >>1 mAP.
6.2.2 Analysis on Multi-Modal Late-Fusion
We study the trade-off between using 2D and 3D RGB detectors, and matching in the 2D image plane and in 3D in Table 3. We further examine the impact of using additional data and study different fusion strategies in Table 4.
How to incorporate RGB information? Table 2 (+ MMF) demonstrates that incorporating RGB information (by matching and filtering 3D LiDAR detections) using a 3D RGB detector greatly improves LT3D. We now evaluate the impact of using 2D RGB-based detectors (e.g., YOLOV7 and DINO) versus 3D RGB-based detectors (e.g., FCOS3D, BEVFormer, PolarFormer), and matching LiDAR detections with 2D RGB detections in 3D versus on the 2D image plane in Table 3. Results show that matching LiDAR detections with 2D RGB detections on the 2D image plane (bottom right) performs best.
Figure 10: Both our MMLF method (columns 1-3) and TransFusion [2] (columns 4 -5) share the same failure cases. In the first and second row, the 2D RGB detector DINO detects the heavily occluded cars but 3D LiDAR detector fails to detect them. As a result, our MMLF misses these cars because it throws away unmatched RGB detections which do not have accurate 3D information. In the third row, we see that, although both the LiDAR and RGB detectors fire on the object (whose ground-truth label is police-officer), the LiDAR detector classifies it as adult and RGB detector classifies it as construction-worker. As a result, the final detection is incorrect w.r.t the predicted label. TransFusion also misclassifies this object as an adult. Refer to Appendix O for more visualizations.
How to match detections from uni-modal detectors? MMF keeps LiDAR detections within a radius of $m$ meters for each RGB detection and removes all the others that are not close to any RGB-based detections. This works well when using a 3D RGB detector, e.g., MMF with FCOS3D, BEVFormer, and PolarFormer improves over the LiDAR-only detector CenterPoint by 2 mAP, as shown in Table 3. Moreover, matching inflated 2D RGB detections in 3D, performs worse than matching 3D RGB detections, achieving marginally lower accuracy than the LiDAR-only baseline.
Notably, projecting LiDAR detections on the 2D image plane and fusing them with 2D RGB detections significantly improves performance for classes with medium and few examples by >>10 mAP. In contrast, projecting 3D RGB detections for matching on the 2D image plane performs worse than matching 2D RGB detections on the 2D image plane, due to the erroneous 3D predictions that affects 2D detection quality.
How to fuse matched detections? Prior to fusion, we first calibrate the scores of LiDAR and RGB detections to ensure that they are comparable. This improves accuracy by 0.7 mAP averaged over all classes (Table 4). With calibration, our probabilistic fusion fuses matched detections to generate the final detections, achieving 0.9 mAP improvements (Table 4).
6.2.3 Analysis of Misclassification
We evaluate CenterPoint [82] and CMT [80] using the hierarchical mAP metric (mAPH), which considers semantic relationships across classes to award partial credit. In safety-critical applications, detecting but misclassifying objects (as a semantically related category) is more desirable than a missed detection (e.g., detecting but misclassifying a child as an adult is preferable to not detecting this child). Table 5 highlights the results of applying our proposed techniques including the group-free detector head, the hierarchy loss, and MMLF using the 2D detector DINO. Results reveal that (1) classes are most often misclassified as their LCA=1 siblings within coarse-grained superclasses, (2) our developed techniques significantly improve mAPH (especially on rare classes), reducing misclassification between sibling classes.
6.2.4 Analysis of Failure Cases
In Fig. 10, we visualize common failure cases of MMLF and compare them with the failure cases of TransFusion [2], an end-to-end trained multi-modal detector. Both TransFusion and our MMLF fail in cases of occlusions (where there is no 3D information) and in cases where the 2D RGB detector misclassifies objects.
Figure 11: Training 2D detectors with more data (e.g., training with nuScenes + nuImages vs. nuScenes only) and using better 2D detectors (e.g., DINO vs. YOLOV7) improve performance on the proxy task of 2D detection and the downstream 3D detection achieved by our multi-modal late-fusion algorithm. See Appendix I for further analysis on the impact of 2D detection quality.
6.2.5 Analysis of Using Strong 2D Detectors
As shown in Fig. 11, simply training better 2D RGB detectors with more data provides a natural pathway for improving LT3D performance. It suggests a strong correlation between 2D detection accuracy and the final 3D LT3D performance. This is an exciting signal, because, as more and more powerful foundational 2D detectors become publicly available [45, 60, 56] , our MMLF framework can utilize these more advanced 2D detectors to further improve LT3D performance.
7 Conclusion
We explore the problem of long-tailed 3D detection (LT3D), detecting objects not only from common classes but also from many rare classes. This problem is motivated by the operational safety of autonomous vehicles (AVs) but has broad robotic applications, e.g., elder-assistive robots [61] that fetch diverse items [23] should address LT3D. To study LT3D, we establish rigorous evaluation protocols that allow for partial credit to better diagnose 3D detectors. We propose several algorithmic innovations to improve LT3D, including a group-free detector head, a hierarchical loss that promotes feature sharing across long-tailed classes, and present a detailed exploration of multi-modal late-fusion for LT3D. We find that 2D RGB detectors are better suited for late-fusion, matching projected 3D LiDAR detections on the 2D image-plane outperforms matching 2D RGB detections inflated to 3D, and score calibration and probabilistic fusion notably improve performance. Our simple multi-modal late-fusion achieves state-of-the-art LT3D performance on the nuScenes benchmark, improving over prior works by 5.9 mAP over all long-tailed distributed classes.
Appendix A Implementation Details
We use the open-source codebase det3d mmcv, mmdetection and mmdetection3d to train FCOS3D, CenterPoint, and TransFusion. We use the first-party implementations for other detectors, following the training schedule proposed by each respective paper. By default, we train the 2D RGB detector with 2D bounding boxes derived from nuScenes’ 3D annotations and additionally train with 2D bounding boxes from nuImages where denoted. Our 2D RGB detectors, YOLOV7 and DINO, are pre-trained on the ImageNet [13] and COCO [36] datasets. We follow the default training recipes for both YOLOV7 and DINO respectively. We describe important implementation details below.
  • Input. We adopt 10-frame aggregation for LiDAR densification when training LiDAR-based detectors on nuScenes and a 5-frame aggregation on Argoverse 2. We assume that we are provided with ego-vehicle poses for prior frames to align all LiDAR sweeps to the current ego-vehicle pose. Since LiDAR returns are sparse, this densification step is essential for accurate 3D detection. By default, we train the 2D RGB detectors on the 2D bounding boxes derived by projecting 3D annotations to the 2D image plane and additionally train with 2D bounding boxes from nuImages where denoted. Our 2D RGB detectors YOLOV7 and DINO are pre-trained on the ImageNet [13] and COCO [36] datasets.
  • Model Architecture. We adopt the architecture in [84] but make an important modification. The original architecture (for the standard nuScenes benchmark) has six heads designed for ten classes; each head has 64 filters. We first adapted this architecture for LT3D using seven heads designed for 18 classes. We then replace these seven heads with a single head consisting of 512 filters shared by all classes.
  • Training Losses. We use the sigmoid focal loss (for recognition) [35] and L1 regression loss (for localization) below. Existing works also use the same losses but only with fine labels; we apply the loss to both coarse and fine labels. Concretely, our loss function for CenterPoint is as follows: $L=L_{HM}+\lambda L_{REG}$, where $L_{HM}=\sum_{i=0}^{C}SigmoidFocalLoss(X_{i},Y_{i})$ and $L_{REG}=|X_{BOX}-Y_{BOX}|$, where $X_{i}$ and $Y_{i}$ are the $i^{th}$ class’ predicted and ground-truth heat maps, while $X_{BOX}$ and $Y_{BOX}$ are the predicted and ground-truth box attributes. Without our hierarchical loss, CC=18. With our hierarchical loss, CC=22 (18 fine grained + 3 coarse + 1 object class). $\lambda$ is set to 0.25. Modifications for other detectors similarly follow.
  • Optimization. We train all LiDAR-only detectors for 20 epoch using an AdamW optimizer and a cyclic learning rate. We adopt a basic set of data augmentations, including global 3D tranformations, flip in BEV, and point shuffling during training. We train our model with 8 RTX 3090 GPUs and a batch size of 1 per GPU. The training noise (from random seed and system scheduling) is << 1% of the accuracy (standard deviation normalized by the mean).
  • Post-processing. We use non-maximum suppression (NMS) on detections within each class to suppress lower-scoring detections. In contrast, existing works apply NMS on all detections across classes, i.e., suppressing detections overlapping other classes’ detections (e.g., a pedestrian detection can suppress other pedestrian and traffic-cone detections).
Appendix B State-of-the-Art Comparison on Argoverse 2
We present results on the large-scale Argoverse 2 (AV2) dataset developed for autonomous vehicle research. AV2 evaluates on 26 classes, which follow a long-tailed distribution. In Fig. 12, we plot the distribution on the left panel and show how we construct a semantic hierarchy on the right panel. Following prior work [46], we train on and evaluate detections up to 50m. As shown in Table 6, our main conclusions from nuScenes still hold for AV2. FCOS3D [76] yields poor performance on all classes, likely due to inaccurate depth estimates. CenterPoint performs considerably better, achieving high accuracy on classes with many examples. Notably, CenterPoint performs better on AV2’s rare classes (30.2 mAP) compared to nuScenes’s rare classes (3.5 mAP), likely because AV2 has more examples per-class in-the-tail. Lastly, our proposed late-fusion approach yields a 8.3 mAPimprovement over CenterPoint and a 3.9 mAP improvement over prior work. These new results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
Few       Medium       Many
Refer to caption Refer to caption
Figure 12: Left: According to the histogram of per-class object counts (on the left), classes in Argoverse 2.0 (AV2) follow a long tailed distribution. Following [37] and nuScenes, we report performance for three groups of classes based on their cardinality (split by dotted lines): Many, Medium, and Few. As AV2 does not provide a class hierarchy, we construct one by referring to the nuScenes hierarchy (cf. Fig. 4 on the right).
Method Modality All Many Medium Few
FCOS3D C 14.6 27.4 17.0 7.8
CenterPoint (CP) L 44.0 77.4 46.9 30.2
MMF(CP, FCOS3D) C + L 48.4 79.0 51.4 35.3
MMLF(CP, DINO) C + L 52.3 89.4 54.2 38.7
Table 6: Comparison with the Argoverse 2 state-of-the-art. We present results on AV2 evaluated at 50m. The RGB-only 3D detector FCOS3D [76] achieves poor performance, likely due to inaccurate depth estimates. In contrast, the LiDAR-only detector CenterPoint (CP) [82] achieves strong performance on all classes. Our multi-modal fusion approach significantly improves over CenterPoint, achieving an 8.3 mAP improvement averaged over all classes. Importantly, late-fusion with 2D detector DINO [88] performs the best by a notable gain. These results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
Appendix C More Visualizations
We present additional visualizations of our multi-modal late-fusion approach in Fig. 13. Our method correctly classifies geometrically similar but semantically different categories like adult versus stroller, bicycle versus personal mobility, child versus adult, and adult versus construction worker. We provide a video of our results in the supplement. Our project page contains a video demo.
Figure 13: We visualize the output of our late-fusion method which combines 3D LiDAR detections from CenterPoint and 2D RGB detections from DINO. In all cases, we find that the 2D RGB detector is able to correct classification errors from the 3D LiDAR detector, improving overall performance. Importantly, score calibration and probabilistic ensembling increases the confidence of the final prediction.
Appendix D Analysis on Transformer 3D Detectors
In the main paper, we evaluate all late-fusion methods by filtering LiDAR detections from CenterPoint with different RGB detectors for fair comparison to [46]. However, our late-fusion approach generalizes to other 3D LiDAR detectors as well. To this end, we evaluate our best-performing late-fusion configuration from the main paper with LiDAR detections from BEVFusion-L [40] and CMT-L [80] in Table 7.
First, we note that BEVFusion-L performs 2.1 mAP better than CenterPoint on LT3D, notably improving on classes with few examples by 6.3%. However, BEVFusion-L performs 4.6 mAP worse than CenterPoint on classes with many examples, suggesting that learning robust features for both common and rare categories is challenging. Surprisingly, CMT-L performs worse than CenterPoint on all classes. We find that filtering BEVFusion-L’s 3D detections with DINO’s 2D detections yields a further 8.2 mAP improvement overall, and a 8.5 mAP improvement for rare classes. This further supports our hypothesis that late-fusion with better 2D detectors can improve LT3D accuracy. In contrast, the end-to-end multi-modal variant of BEVFusion only improves over the LiDAR-only baseline by 3 mAP overall, and performs 5.2 mAP worse than our late-fusion approach. Similarly, filtering CMT-L’s 3D detections with DINO’s 2D detections yields a 15.9 mAP improvement overall, with a 19.1 mAP improvement for rare classes. Fusing CMT-L with DINO performs marginally better than CenterPoint with DINO.
We posit that the relative improvement from late-fusion can be attributed more generally to LiDAR-based detectors achieving high recall and low precision rather than the specific network architecture. Our late-fusion approach will be less effective for LiDAR-based detectors that achieve high precision and low recall since our method does not add new 3D detections (by lifting 2D detections), but rather only filters existing ones.
Method All Many Medium Few
CenterPoint w/ Hier. 40.4 77.1 45.1 4.3
MMLF(CenterPoint, DINO) 51.4 77.9 59.4 20.0
BEVFusion-L 42.5 72.5 48.0 10.6
MMLF(BEVFusion-L, DINO) 50.7 76.8 59.1 19.1
CMT-L 34.7 73.4 35.9 1.1
MMLF(CMT-L, DINO) 51.6 79.1 58.8 20.2
BEVFusion (RGB + LiDAR) 45.5 75.5 52.0 12.8
CMT (RGB + LiDAR) 44.4 79.9 53.0 4.8
Table 7: Transformer-based 3D LiDAR detector. We evaluate the late-fusion performance of BEVFusion-L and CMT-L (LiDAR-only) with DINO, and find that our late fusion strategy improves over BEVFusion-L by 8.2 mAP overall and CMT-L by 15.9 mAP. This suggests that our late-fusion approach generalizes across different 3D detector architectures.
Appendix E Inference Runtime
FCOS3D BEVFormer CenterPoint (CP) TransFusion DeepInteraction MMLF(CP, DINO)
[76] [43] [82] [2] [79] (Ours)
Time (ms) 89 327 323 367 590 323
mAP 20.9 27.3 39.2 39.8 43.7 51.4
Table 8: Comparison of inference time. Our multi-modal late-fusion approach runs faster than existing methods including end-to-end multimodal detectors such as DeepInteraction. Note that although our approach runs two uni-modal detectors, we run them in parallel so that the detection stage maintains the same efficiency as CenterPoint (CP). Moreover, our late-fusion strategy introduces negligible computation in inference. We copy 3D detection performance (in mAP) from the main paper for reference. Results demonstrate that our simple multi-modal late-fusion approach resoundingly outperforms sophisticated state-of-the-art methods (e.g., DeepInteraction and TransFusion) with significantly less inference time. Nevertheless, all methods require implementation optimization towards real-time processing for real-world application.
Table 8 compares the inference runtime of our method against prior works on a single GPU. Although our multi-modal late-fusion (MMLF) methods such as MMLF(CenterPoint, DINO) run two uni-modal detectors, we run them in parallel. Therefore, our final detector maintains the same runtime as a typical uni-modal detector and is often faster than more sophisticated multimodal detectors like BEVFormer [43], TransFusion [2] and DeepInteraction [79]. However, running two detectors in parallel will require more memory usage. Importantly, late-fusion methods including MMLF(CenterPoint, DINO) and [46] have negligible overhead when fusing uni-modal detections, hence their runtime is the same as CenterPoint, which serves as the speed bottleneck. This work uses research-level code and can be further optimized for deployment on autonomous vehicles.
Appendix F Analysis on Class Grouping
CenterPoint Multi-Head Car Ped. Barrier TC Truck Bus Trailer CV MC Bicyc.
Original 87.7 87.7 70.7 74.0 63.6 72.7 45.1 26.3 64.7 47.9
89.1 88.4 70.8 74.3 64.8 72.9 42.0 25.7 65.9 53.6
for LT3D 82.4 62.0 60.1 49.4 55.7 28.9 19.7 48.9 33.6
88.1 72.4 72.7 62.7 70.8 40.2 24.5 62.8 48.5
Adult $P$ CW Debris Child Stroller PO EV PM All
Original 64.0
64.8
for LT3D 81.2 21.7 14.2 1.1 0.1 0.1 1.3 0.1 0.1 31.2
86.3 32.7 22.2 4.3 0.1 4.3 1.8 10.3 0.1 39.2
Table 9: Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D  benchmarks. We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach. Note, TC is traffic-cone, CV is construction vehicle, MC is motorcycle, $P$ is pushable-pullable, CW is construction-worker, and PO is police-officer. We highlight Medium and Few classes in blue.
Many contemporary networks use a multi-head architecture that groups classes of similar size and shape to facilitate efficient feature sharing. For example, CenterPoint groups pedestrian and traffic-cone since these objects are both tall and skinny. We study the impact of grouping for both the standard and LT3D problem setups. We define the groups used for this study below. Each group is enclosed in curly braces. Our group-free head includes all classes into a single group.
  • Original: {Car}, {Truck, Construction Vehicle}, {Bus, Trailer}, {Barrier}, {Motorcycle, Bicycle}, {Pedestrian, Traffic Cone}
  • LT3D: {Car}, {Truck, Construction Vehicle}, {Bus, Trailer}, {Barrier}, {Motorcycle, Bicycle}, {Adult, Child, Construction Worker, Police Officer, Traffic Cone}, {Pushable Pullable, Debris, Stroller, Personal Mobility, Emergency Vehicle}
We use the class groups proposed by prior works [84, 82] for the standard benchmark and adapt this grouping for LT3D. Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D benchmarks, as shown in Table 9. We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach. The group-free approach simplifies architecture design, while also providing competitive performance.
Two insights allow us to train the group-free architecture. First, we make the group-free head proportionally larger to train more classes. The standard grouping setup contains 6 heads, each with 64 convolutional filters. Scaling up to the nearest power of two, our group-free head has 512 convolutional filters. Second, we do not perform between-class NMS. The standard setup performs NMS between classes in each group (e.g., since pedestrians and traffic cones are tall and skinny, the model should only predict that an object is either a traffic cone or a pedestrian). However, performing NMS between classes requires that confidence scores are calibrated, which is not the case. Moreover, for LT3D, score calibration becomes more important for rare classes as these classes have lower confidence scores than common classes on average, meaning that common objects will likely suppress rare objects within the same group. Our solution is to only perform within-class NMS, which is standard for 2D detectors [58].
Appendix G Analysis of Hierarchical Training
Method Hier. Many Medium Few All
CenterPoint (w/o Hier.) n/a 76.4 43.1 3.5 39.2
CenterPoint w/ Hier. (a) 77.1 45.1 4.3 40.4
(b) 76.4 45.0 5.3 40.5
(c) 76.5 45.2 5.2 40.6
(d) 74.5 43.5 5.6 39.5
Table 10: Different variants achieve similar performance. We note that other methods do improve accuracy in the tail by sacrificing performance in the head, suggesting that hybrid approaches that apply different techniques for head-vs-tail classes may further improve accuracy. Unlike [42, 74] which requires a strict label hierarchy, our approach is not limited to a hierarchy.
Classic methods train a hierarchical softmax (in contrast to our simple approach of sigmoid focal loss with both fine and coarse classes), where one multiplies the class probabilities of the hierarchical predictions during training and inference [74]. We implemented such an approach, but found the training did not converge. Interestingly, [74] shows such a hierarchical softmax loss has little impact on long-tailed object detection (in 2D images), which is one reason they have not been historically adopted. Instead, we found better results using the method from [42] (a winning 2D object detection system on the LVIS [19] benchmark) which multiples class probabilities of predictions (e.g. $P_{CAR}=P_{OBJ}*P_{CAR}$) at test-time, even when such predictions are not trained with a hierarchical softmax. We tested three variants and compared it to our approach (which recall, uses only fine-grained class probabilities at inference). Table 10 compares their performance for LT3D.
Class name w/ YOLOV7 w/ DINO
Personal Mobility 12.5 17.3
Emergency Vehicle 10.4 16.2
Police Officer 10.7 20.9
Stroller 22.1 32.3
Child 6.4 7.6
Debris 5.2 5.9
Construction Worker 40.4 54.6
Bicycle 63.1 68.2
Motorcycle 72.0 74.1
Construction Vehicle 31.7 33.6
Bus 75.3 75.9
Pushable Pullable 46.2 54.2
Trailer 40.8 43.3
Truck 58.5 60.1
Traffic Cone 79.9 80.5
Barrier 74.2 74.3
Adult 85.8 86.0
Car 86.9 86.1
Few 11.2 16.7
Medium 52.8 57.7
Many 77.1 77.4
All 45.7 49.5
Table 11: Using a stronger 2D detector improves 3D detection by MMLF. We compare using DINO [88] vs. YOLOv7 [70] as the 2D detector, applied along with the 3D LiDAR detector CenterPoint [82] in our MMLF. Results demonstrate that using DINO, a stronger detector than YOLOv7, yields significantly better 3D detection performance on all classes, particularly on medium and few classes. Refer to Fig. 1 for the distribution of these classes.
  1. (a)
    Ours (e.g., Fine-class score only)
  2. (b)
    Object score * Fine-class score ([42], e.g. $P_{CAR}=P_{OBJ}*P_{CAR}$)
  3. (c)
    Coarse score * Fine-class score (Variant-1 of [42], e.g. $P_{CAR}=P_{VEHICLE}*P_{CAR}$)
  4. (d)
    Object score * Coarse score * Fine-class score (Variant-2 of [42], e.g. $P_{CAR}=P_{OBJ}*P_{VEHICLE}*P_{CAR}$)
Unlike [42, 74] which require a strict label hierarchy, our approach is not limited to a hierarchy. We find that other hierarchical methods improve accuracy in the tail by sacrificing performance in the head, suggesting that hybrid approaches that apply different techniques for head-vs-tail classes may further improve accuracy.
Appendix H Per-Class Breakdown Results
Class name BevFormer Deepinteraction CMT MMLF(CenterPoint, DINO)
[43] [79] [80] (Ours)
Personal Mobility 0.7 2.5 4.1 19.3
Emergency Vehicle 0.2 1.5 6.9 21.4
Police Officer 0.9 0.0 7.3 25.6
Stroller 1.1 30.9 0.9 37.7
Child 3.9 6.0 4.7 8.8
Debris 1.9 6.8 4.6 6.8
Construction Worker 21.5 30.7 34.4 55.9
Bicycle 41.9 63.3 66.7 70.0
Motorcycle 47.0 76.2 76.3 75.9
Construction Vehicle 16.9 35.3 36.9 35.3
Bus 43.0 75.9 76.4 76.1
pushable Pullable 29.1 30.8 34.1 58.1
Trailer 21.4 45.4 46.5 44.6
Truck 37.1 63.2 65.2 60.6
Traffic Cone 59.9 71.8 80.1 81.5
Barrier 53.2 75.3 77.9 74.6
Adult 50.3 85.9 87.7 86.2
Car 60.8 84.9 88.6 86.3
Few 1.4 7.9 4.8 20.0
Medium 31.6 51.1 53.0 59.5
Many 52.3 76.2 79.9 77.9
All 27.3 43.7 44.4 51.4
Table 12: Per-class performance on nuScenes. Our approach MMLF fuses detections from the LiDAR-based 3D detector CenterPoint and the RGB-based 2D detector DINO, achieving the highest per-class AP on 10 out of 18 classes. It is worth noting that, compared to CMT, our approach improves construction worker by 21.5 AP, stroller by 36.8 AP, and pushable-pullable by 24.0 AP.
We compare recent multi-modal methods w.r.t per-class AP in Table 12. All multi-modal methods perform similarly on common classes but considerably worse on Medium and Few classes, highlighting the need for further investigation of LT3D by the research community. Notably, our late-fusion approach achieves 20.0 higher AP on pushable-pullable and 6.0 higher AP on stroller than prior work. In general, our late-fusion approach yields considerable improvement on classes with medium and few examples. Despite significant improvements in rare class detection accuracy, our approach detects child with a low 8 mAP, which is still much better than compared methods. We posit that it is difficult to differentiate child from adult due to perspective geometry, since a small child close to the camera looks similar to a tall adult far away from the camera.
Appendix I Analysis of 2D Detection Performance
Method All Many Medium Few
FCOS3D (3D RGB Detector) 18.3 36.0 21.1 0.2
BEVFormer (3D RGB Detector) 23.0 40.8 28.2 2.1
PolarFormer (3D RGB Detector) 20.7 37.5 25.2 1.6
YOLOV7 (2D RGB Detector) 37.5 63.5 45.0 7.1
DINO (2D RGB Detector) 44.3 67.8 51.9 15.9
Table 13: Evaluating 2D detection performance. We find that simply projecting 3D detections from RGB-only 3D detectors to 2D image plane yields considerably lower 2D detection mAP across all classes. However, 2D RGB detectors achieve higher performance. Intuitively, 2D RGB detectors achieve better classification performance and predict tighter fitting bounding boxes than those derived by projecting 3D detections to the image plane. This explains why fusing detections from 2D RGB detectors yields better LT3D performance than detections from 3D RGB detectors. All detectors in this table are only trained on the nuScenes train-set.
Table 13 shows that using the 2D detector DINO outperforms state-of-the-art 3D RGB detectors (e.g., BEVFormer) for 2D detection on the nuScenes val-set. Importantly, DINO performs significantly better than BEVFormer (15.9 vs. 2.1 mAP) on rare classes. Furthermore, Table 11 compares using DINO [88] vs. YOLOv7 [70] as the 2D detector, applied along with the 3D LiDAR detector CenterPoint [82] in our MMLF. Results show that using DINO, a stronger detector than YOLOv7, yields significantly better 3D detection performance on all classes, particularly on medium and few classes.
Appendix J Analysis of AP vs. NDS Results
We report nuScenes Detection Score (NDS) and mAP results in Table 14. We find that all prior methods follow the same ranking on both metrics. This is unsurprising because NDS is computed as a weighted sum of mAP and other true positive metrics, where mAP is weighted five times greater than other components.
FCOS3D BEVFormer CP TransFusion DeepInteraction CMT MMLF(CP, DINO)
[76] [43] [82] [2] [79] [80] (Ours)
mAP 20.9 27.3 39.2 39.8 43.7 44.4 51.4
NDS 30.4 38.8 54.9 53.9 54.4 55.9 60.4
Table 14: AP vs. NDS results. We compare methods reported in the main paper w.r.t nuScenes Detection Score (NDS) and mAP. We find that the methods follow the same rankings on both metrics.
Appendix K Standard nuScenes Benchmark Results
Method Car Truck Trailer Bus CV Bicy. MC Ped. TC Barrier mAP NDS
CenterPoint [82] 87.7 61.6 43.7 73.4 28.4 49.5 65.9 87.1 75.7 71.3 66.6 64.4
MMLF(CenterPoint, DINO) 86.2 60.1 44.6 76.1 35.3 70.0 75.9 86.6 81.5 74.6 69.2 71.3
CMT [80] 88.6 65.1 46.5 76.4 36.9 66.7 76.2 88.1 80.1 77.9 70.3 72.9
MMLF(CMT, DINO) 88.3 64.8 46.2 75.8 36.6 69.9 76.4 87.7 81.5 77.5 70.5 73.1
Table 15: 3D detection performance on the standard nuScenes benchmark. We evaluate the impact of late-fusion for the standard nuScenes benchmark. Although our work focuses on improving detection accuracy in-the-tail, we find that our proposed approach can still improve detection accuracy for some common classes in the standard benchmark. Notably, fusing CenterPoint [82] with DINO [88] improves detection accuracy by 2.6 mAP and 6.9 NDS. However, fusing CMT [80] (a multi-modal 3D detector) with DINO [88] increases detection accuracy by 0.2 mAP, only improving on bicycle and traffic-cone. We posit that CMT already effectively uses multi-modal training data for the (common) classes of the standard nuScenes benchmark, yielding diminishing returns with late-fusion. Note that CV stands for construction vehicle, MC represents motorcycle, and TC denotes traffic cone.
In Table 15, we evaluate our late-fusion method on the standard 10 classes of the nuScenes benchmark [5]. Although our late-fusion approach primarily improves performance for classes with medium and few examples per class, we find that the late-fusion of CenterPoint [82] and DINO [88] yields a 2.6 and 6.9 performance increase for mAP and NDS on the standard nuScenes benchmark, respectively. Notably, we see significant perofrmance improvements for bicycle and motorcycle, likely because our LiDAR-only detector misclassifies these geometrically similar, but semantically different categories. Next We apply our late-fusion to CMT [80], a state-of-the-art multi-modal 3D detector, but find that our late-fusion approach only provides marginal performance improvements. Late-fusion of CMT [80] and DINO [88] only marginally improves bicycle and traffic cone performance. This suggests that CMT is already able to effectively utilize multi-modal information.
Appendix L Impact of Occlusion and Range
We report average recall at four visibility levels in Table 16. Results show that detecting occluded and partially occluded objects is significantly more challenging than detecting un-occluded objects, likely because the visible regions of partially occluded objects can be easily mistaken as part of the background and hence lead to more false positive detections. Interestingly, the LiDAR-based 2D detector CenterPoint achieves the highest recall across all occlusion levels.
We also examine detection accuracy across three distance ranges in Table 17 and observe that detecting objects closer to the ego-vehicle is easier due to the higher LiDAR point density. While CenterPoint’s performance declines significantly beyond 10m, CMT and MMLF demonstrate greater robustness, suggesting that both feature-fusion and late-fusion effectively mitigate the impact of LiDAR sparsity at range. Furthermore, our multi-modal fusion approach significantly improves CenterPoint’s detection accuracy in the 10 – 20m and 20 – 30m ranges by reducing false positives, increasing AP from 47.9 to 63.2 and from 36.0 to 44.5, respectively.
Method Modality 0-40% 40-60% 60-80% 80-100%
CenterPoint (CP) L 66.0 78.7 79.5 80.1
CMT C + L 56.5 73.5 75.0 75.8
MMLF(CP, DINO) C + L 49.1 67.3 68.3 78.7
Table 16: Average recall for occluded targets at different visibility levels. We evaluate average recall at four visibility levels (defined by nuScenes). Perhaps unsurprisingly, detection accuracy drops as objects become more occluded. Note that visibility attributes are not predicted by our model and are manually annotated for the ground truth, making it difficult to measure precision.
Method Modality 0-10m 10-20m 20-30m Full Range
CenterPoint (CP) L 60.3 47.9 36.0 40.4
CMT C + L 56.9 51.5 39.8 44.4
MMLF(CP, DINO) C + L 63.2 63.2 44.5 51.4
Table 17: Detection accuracy at different distance thresholds. We evaluate detection accuracy at three distance ranges. Detecting objects closer to the ego-vehicle is easier (due to higher LiDAR point density) than detecting objects further away.
Appendix M Analysis of Recall
As shown in Table 18, the LiDAR-only 3D detector CenterPoint achieves higher recall than CMT (multi-modal) for the Pedestrian (LCA=1), Movable (LCA=1), and Object (LCA=2) superclasses. The lower recall of CMT can be attributed to its limited number of object queries (N=300) per frame compared to the (nearly) unbounded number of predictions from anchor-free detectors like CenterPoint. Importantly, the high recall of CenterPoint is essential for our late-fusion approaches (cf. Fig. 6).
Method Modality Vehicle Pedestrian Movable Object
CenterPoint L 85.5 95.5 90.7 89.3
CMT C + L 86.4 91.2 87.2 87.9
Table 18: Average recall between detectors. CenterPoint (LiDAR-only) achieves higher recall than CMT for the Pedestrian (LCA=1), Movable (LCA=1), and Object (LCA=2) superclasses.
Appendix N Impact of Calibration Noise
Translation Noise $N_{T}$ 0.01m 0.02m 0.03m 0.04m
CMT 44.4 44.4 44.4 44.4
MMLF(CP, DINO) 51.4 51.5 51.5 51.5
Rotation Noise $N_{R}$ 1 2 3 4
CMT 44.3 43.1 42.7 41.1
MMLF(CP, DINO) 48.7 47.8 46.1 45.6
Translation Noise $N_{T}$ 0.01m 0.02m 0.03m 0.04m
Rotation Noise $N_{R}$ 1 2 3 4
CMT 44.3 43.8 43.1 42.2
MMLF(CP, DINO) 47.8 46.6 45.1 44.6
Table 19: Impact of calibration noise. We apply a small amount of additive Gaussian noise to the translation and rotation calibration matrices to measure the impact of imprecise calibration on late fusion. While translation noise has little effect on detection accuracy, rotation noise significantly impacts performance. As expected, combining both translation and rotation noise degrades accuracy further.
Following prior literature, we benchmark on datasets (nuScenes and Argoverse 2) collected in-the-wild, which already contain realistic calibration errors. However, we study the impact of introducing more aggressive calibration errors in Table 19 by applying synthetic perturbations to the camera-to-LiDAR translation and rotation matrices. We apply additive Gaussian noise $N_{T}$ to each of the $X$, $Y$ and $Z$ components of the translation vector and apply additive noise $N_{R}$ to the yaw, pitch and roll of the rotation matrix. We find that translation noise does not significantly affect detection accuracy. In contrast, rotation noise has a much larger impact on accuracy. Unsurprisingly, combining translation and rotation noise yields significantly degraded accuracy. Such calibration errors affect all (early, mid, and late) fusion methods since they require calibrated extrinsics [47]. We find that our late-fusion approach is less robust to noise than end-to-end fusion methods. However, our approach still achieves higher accuracy than CMT.
Appendix O 3D Visualization of Select Figures
In Fig. 14 and 15, we visualize the LT3D detection results in a 3D virtual scene. The results are also visualized in the main paper (corresponding to Fig. 9 and 10, respectively) in the BEV scene.
MMLF(CenterPoint, DINO)           CenterPoint
Refer to caption Refer to caption
Figure 14: Three examples demonstrate how our multi-modal late-fusion (MMLF) approach improves LT3D by ensembling 2D RGB detections (from DINO [88]) and 3D LiDAR detections (from CenterPoint [82]). In all examples, MMLF correctly relabels detections which are geometrically similar (w.r.t size and shape) in LiDAR but are visually distinct in RGB, such as bus-vs-truck, adult-vs-stroller, and adult-vs-child. This figure complements Fig. 9 in the main paper.
MMLF(CenterPoint, DINO)           TransFusion
Figure 15: Both our MMLF method (column 1) and TransFusion [2] (column 2) share the same failure cases. In the first and second row, the 2D RGB detector DINO detects the heavily occluded cars but 3D LiDAR detector fails to detect them. As a result, our MMLF misses these cars because it throws away unmatched RGB detections which do not have accurate 3D information. In the third row, we see that, although both the LiDAR and RGB detectors fire on the object (whose ground-truth label is police-officer), the LiDAR detector classifies it as adult and RGB detector classifies it as construction-worker. As a result, the final detection is incorrect w.r.t the predicted label. TransFusion also misclassifies this object as an adult. This figure complements Fig. 10 in the main paper.